OpenStudy (anonymous):

Differentiate. f(θ) = sec θ/3 + sec θ

OpenStudy (anonymous):

\[f(t) = \frac{ \sec(t) }{ 3 } + \sec(t)\]

OpenStudy (anonymous):

this is the question ?

OpenStudy (anonymous):

or it's 3+sec(t) in the denominator ?

OpenStudy (anonymous):

yeaaa

OpenStudy (anonymous):

in the denominator

OpenStudy (anonymous):

\[f(t)=\frac{ \sec(t) }{ 3 + \sec(t) }\]

OpenStudy (anonymous):

like this ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

In this case, use quotient rule.

OpenStudy (anonymous):

grr mistake :| all over again

OpenStudy (anonymous):

\[f'(t) = \frac{ \sec(t)' * (3+\sec(t)) - (3+\sec(t))' * \sec(t) }{ (3+\sec(t))^2 }\]

OpenStudy (anonymous):

\[f'(t) = \frac{ \tan(t)\sec(t) * (3+\sec(t)) - \tan(t)\sec(t) * \sec(t) }{ (3+\sec(t))^2 }\]

OpenStudy (anonymous):

\[f'(t) = \frac{ 3\tan(t)\sec(t) }{ (3+\sec(t))^2 }\]

OpenStudy (anonymous):

nothing much more to do ..

OpenStudy (anonymous):

so you expand and simplify?

OpenStudy (anonymous):

the final answer ? you can expand the (3+sec(t))^2 but it's not really important ..

OpenStudy (anonymous):

i used the quotient rule in order to find the derivative as you can see the procedure

OpenStudy (anonymous):

okaay thank you :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

i got the aanswer wrong

OpenStudy (anonymous):

i'm doing this online quiz and wen i put in the answer it was wrong